# What is the Curvature of 2D Euclidean Quantum Gravity?

### v1

@@ 5.2.3 Finite-size effects
To address this issue, we start by implementing a relatively crude method where we eliminate the volumes deemed too small to evade significant finite-size effects, while keeping the choice $\delta_{\mathrm{max}}\!=\!15$ fixed. In the absence of any unique or sharp criteria, we work with the requirements that (i) configurations with a maximal diameter 45 should “fit inside” the sphere-like geometry in the sense that $3\delta_{\mathrm{max}}\!\equiv\!45\!\lesssim\!r_{\mathrm{max}}$, where the average maximal linear extension $r_{\mathrm{max}}$ for a given ensemble and volume $N_{2}$ is read off from Fig. 9, and (ii) the data points of the included volumes should lie approximately on a straight line in the log-log plots like those shown in Figs. 12 and 15. This leads to the exclusion of the effective curvature radii $\rho_{\mathrm{eff}}$ for the system sizes $N_{2}\!<\!60$k for all three ensembles, corresponding to the three lowest-lying data points with $\log N_{2}\!<\!11.0$ in the log-log plots. The resulting power-law fits for the particular case $D\!=\!4$ are shown in Fig. 16. Unsurprisingly, the reduced volume range leads to a much better quality of the linear fits, which also exhibit clearly that the omitted data points are associated with a flatter slope. The best sphere fitting dimensions $\cal D$ extracted from the scaling ansatz (25) for all ensembles are collected in Table 4. A significant change has occurred in comparison with the fits that included all volumes: for all three ensembles, the dimensions $\cal D$ and $D$ now agree best for $D\!=\!4$, at the expense of somewhat larger error bars.
@@
The resulting fits for $\log(\rho_{\mathrm{eff}})$ as a function of $\log(N_{2})$, for $D\!=\!4$ and all three ensembles, are presented in Fig. 17. It is gratifying to see that our refined methodology appears to have eliminated successfully the finite-size effects for the smaller volumes, in the sense that all data points now lie beautifully along straight lines. The results for the best sphere fitting dimension $\cal D$ extracted from the fits for all ensembles are presented in Table 5. Compared with our previous method, which omitted all data points for volumes $N_{2}\!<\!60$k, the error bars have been approximately halved. As a further improvement, for $D\!=\!4$, the values for $D$ and $\cal D$ now coincide within error bars (for ${\cal T}_{c}$ and ${\cal T}_{r}$) or lie just outside the error bar (for ${\cal T}_{m}$). The excellent quality of this agreement leads us to conclude that the local curvature properties of 2D Euclidean quantum gravity resemble closest those of a continuum sphere of dimension four, contrary to the earlier finding of five as the preferred dimension [9].
@@
5.2.4 Summary of improvements
Let us review and summarize the findings of our improved treatment of the curvature profile. We started out in Sec. 5.2.1 by replicating the analysis of reference [9] for the combinatorial ensemble, with the same choices $\delta_{\mathrm{min}}\!=\!5$ and $\delta_{\mathrm{max}}\!=\!15$. Despite some discrepancies in the data we reached the same conclusion as [9], namely that the best match between the dimension $D$ of the comparison sphere in the continuum and the dimension $\cal D$ obtained from the scaling of the effective radius $\rho_{\mathrm{eff}}$ with the lattice volume is realized for $D\!=\!5$. Performing the same analysis for the two degenerate ensembles at first sight appeared to confirm this result, but the poor quality of the log-log fits, illustrated by Fig. 12, alerted us to the possible presence of discretization artefacts and finite-size effects. Our subsequent investigation of these effects in Sec. 5.2.2 led us to using a larger threshold $\delta_{\mathrm{min}}\!=\!7$ for the combinatorial ensemble, to evade short-distance lattice artefacts still present at $\delta_{\mathrm{min}}\!=\!5$, and the use of $\delta_{\mathrm{min}}\!=\!6$ for the two other ensembles. Implementing $\delta_{\mathrm{min}}\!=\!7$ in the combinatorial ensemble revealed finite-size effects (cf. Fig. 15), whose absence up to that moment had been conspicuous, given the presence of these effects in the measurements for the degenerate ensembles.
Sec. 5.2.3 dealt with the systematic elimination of these finite-size effects. Our initial method consisted in discarding all data pairs $(N_{2},\rho_{\mathrm{eff}}(N_{2}))$ for lattice volumes $N_{2}\!<\!60$k. The disadvantage of this prescription is that more data than strictly necessary are eliminated, leading to larger statistical errors. We remedied this by adopting an alternative method, where we only discarded selected, individual data points $(\delta,\langle\bar{d}_{\mathrm{av}}(\delta)/\delta\rangle_{N_{2}})$ from the fits to the spherical curvature profiles, namely those with $\delta$ larger than a $N_{2}$-dependent threshold value $\delta_{\mathrm{max}}(N_{2})$. This resulted in new, corrected data pairs $(N_{2},\rho_{\mathrm{eff}}(N_{2}))$ for the volumes $N_{2}\!\leq\!60$k.
@@ 5.2.4 Summary of improvements
Both methods led to the same conclusion: the sphere dimension $D$ and the sphere fitting dimension $\cal D$ agree best for $D\!=\!4$, for all ensembles. Particularly notable was the outcome of the second, improved method, which makes optimal use of the collected data, where the logarithmic fits are of a very good quality and the agreement is approximately within one error bar (cf. Table 5). It provides strong evidence that the averaged local curvature of 2D Euclidean quantum gravity, as captured by the curvature profile, matches closely that of a continuum four-sphere, and that this result is independent of the chosen ensemble of triangulations, demonstrating universality. This both corrects and improves on previous results in [9].
To further illustrate the progress we have made, let us examine the “curve collapse” of Fig. 6 in [9] for the combinatorial ensemble, which collects the curvature profiles $\langle\bar{d}_{\mathrm{av}}(\delta)/\delta\rangle_{N_{2}}$ as a function of the rescaled distances $\delta/\rho_{\mathrm{eff}}$ for all volumes $N_{2}$. If they fall on a common curve, it indicates the presence of finite-size scaling, with a well-defined infinite-volume limit. Fig. 18 shows the corresponding graphs before and after the improvements we have described in this Section, together with continuum curves for the curvature profile of a round four-sphere with a best vertical shift. We observe that in the plot on the right the overlap among data for different volumes and their closeness to the continuum curve is much improved, for the subset of points considered, lending further support to our conclusion.
@@ 6 Summary and outlook
We began by measuring average sphere distances according to the specifications of [9], i.e. in the ensemble ${\cal T}_{c}$ of combinatorial triangulations and using data points in the interval $\delta\!\in\![5,15]$. At first glance this appeared to confirm the previous conclusion that the preferred sphere dimension is $D\!=\!5$, until we performed the same analysis for the two degenerate ensembles ${\cal T}_{r}$ and ${\cal T}_{m}$, which strongly suggested the presence of lattice artefacts. As summarized in Sec. 5.2.4, re-examining this issue in all three ensembles and carefully selecting data points deemed unaffected by such artefacts resulted in fits of very good quality (cf. Fig. 17), an improved finite-size scaling (illustrated by Fig. 18) and a new and different conclusion: it is not for $D\!=\!5$ but for $D\!=\!4$ that the two dimensions $D$ and $\cal D$ coincide, and they do so within error bars.

### v2

@@ 5.2.3 Finite-size effects
To address this issue, we start by implementing a relatively crude method where we eliminate the volumes deemed too small to evade significant finite-size effects, while keeping the choice $\delta_{\mathrm{max}}\!=\!15$ fixed.
@@
In the absence of any unique or sharp criteria, we work with the requirements that (i) configurations with a maximal diameter 45 should “fit inside” the sphere-like geometry in the sense that $3\delta_{\mathrm{max}}\!\equiv\!45\!\lesssim\!r_{\mathrm{max}}$, where the average maximal linear extension $r_{\mathrm{max}}$ for a given ensemble and volume $N_{2}$ is read off from Fig. 4, and (ii) the data points of the included volumes should lie approximately on a straight line in the log-log plots like those shown in Figs. 7 and 10. This leads to the exclusion of the effective curvature radii $\rho_{\mathrm{eff}}$ for the system sizes $N_{2}\!<\!60$k for all three ensembles, corresponding to the three lowest-lying data points with $\log N_{2}\!<\!11.0$ in the log-log plots. The resulting power-law fits for the particular case $D\!=\!4$ are shown in Fig. 11. Unsurprisingly, the reduced volume range leads to a much better quality of the linear fits, which also exhibit clearly that the omitted data points are associated with a flatter slope. The best sphere fitting dimensions $\cal D$ extracted from the scaling ansatz (24) for all ensembles are collected in Table 4. A significant change has occurred in comparison with the fits that included all volumes: for all three ensembles, the dimensions $\cal D$ and $D$ now agree best for $D\!=\!4$, at the expense of somewhat larger error bars.
@@
The resulting fits for $\log(\rho_{\mathrm{eff}})$ as a function of $\log(N_{2})$, for $D\!=\!4$ and all three ensembles, are presented in Fig. 12. It is gratifying to see that our refined methodology appears to have eliminated successfully the finite-size effects for the smaller volumes, in the sense that all data points now lie beautifully along straight lines. The results for the best sphere fitting dimension $\cal D$ extracted from the fits for all ensembles are presented in Table 5.
@@
Compared with our previous method, which omitted all data points for volumes $N_{2}\!<\!60$k, the error bars have been approximately halved. As a further improvement, for $D\!=\!4$, the values for $D$ and $\cal D$ now coincide within error bars (for ${\cal T}_{c}$ and ${\cal T}_{r}$) or lie just outside the error bar (for ${\cal T}_{m}$). The excellent quality of this agreement leads us to conclude that the local curvature properties of 2D Euclidean quantum gravity resemble closest those of a continuum sphere of dimension four, contrary to the earlier finding of five as the preferred dimension [9].
To further illustrate the progress we have made relative to previous results, let us examine the “curve collapse” of Fig. 6 in [9] for the combinatorial ensemble, which collects the curvature profiles $\langle\bar{d}_{\mathrm{av}}(\delta)/\delta\rangle_{N_{2}}$ as a function of the rescaled distances $\delta/\rho_{\mathrm{eff}}$ for all volumes $N_{2}$. If they fall on a common curve, it indicates the presence of finite-size scaling, with a well-defined infinite-volume limit. Fig. 13 shows the corresponding graphs before and after the improvements we have described in this Section, together with continuum curves for the curvature profile of a round four-sphere with a best vertical shift. We observe that in the plot on the right the overlap among data for different volumes and their closeness to the continuum curve is much improved, for the subset of points considered, lending further support to our conclusion.
@@ 6 Summary and outlook
We began by measuring average sphere distances according to the specifications of [9], i.e. in the ensemble ${\cal T}_{c}$ of combinatorial triangulations and using data points in the interval $\delta\!\in\![5,15]$. At first glance this appeared to confirm the previous conclusion that the preferred sphere dimension is $D\!=\!5$, until we performed the same analysis for the two degenerate ensembles ${\cal T}_{r}$ and ${\cal T}_{m}$, which strongly suggested the presence of lattice artefacts. Re-examining this issue in all three ensembles and carefully selecting data points deemed unaffected by such artefacts resulted in fits of very good quality (cf. Fig. 12), an improved finite-size scaling (illustrated by Fig. 13) and a new and different conclusion: it is not for $D\!=\!5$ but for $D\!=\!4$ that the two dimensions $D$ and $\cal D$ coincide, and they do so within error bars.
@@ Acknowledgments
Appendix
In this appendix, we define the three ensembles in terms of their geometric properties as (generalized) triangulations, which is how they appear in the numerical investigations. We describe in more detail the geometric irregularities of triangulations associated with tadpoles and self-energy subgraphs in the dual graph picture, and how these different characterizations relate to each other. Although bits and pieces are covered and referred to in the literature (see e.g. [34, 35, 38]), we have not found a place where all of these aspects are discussed together.
Our starting point are finite triangulations with the topology of a two-sphere $S^{(2)}$, whose elementary building blocks are flat equilateral triangles (two-simplices $\sigma^{(2)}$) of edge length $a$, the already mentioned UV-cutoff. $D$-simplices $\sigma^{(D)}$ of lower dimension $D$ contained in the triangles are their edges or links (one-simplices $\sigma^{(1)}$) and vertices (zero-simplices $\sigma^{(0)}$). The ensembles differ by how the triangles can be glued together pairwise along their edges to obtain a piecewise flat geometry $T$.
In what follows we will need a few definitions regarding simplicial structures [34, 36]. A face of a simplex $\sigma$ is any simplex whose vertices are a subset of the vertices of $\sigma$. The star st$(\sigma)$ of a simplex $\sigma$ in a triangulation $T$ is the union of all simplices in $T$ that have $\sigma$ as a face. Lastly, the link lk$(\sigma)$ of a simplex $\sigma$ is the union of all simplices $\tau\!\in\!\mathrm{st}(\sigma)$ which have vanishing intersection with $\sigma$, $\tau\cap\sigma\!=\!\emptyset$.
Fig. 14 shows an example of the star and link of a zero-simplex $\sigma^{(0)}$. [^19] The notion of “link” introduced here should not be confused with what elsewhere we use as a synonym for “edge”. A triangulation $T$ of $S^{(2)}$ is combinatorial, $T\!\in\!{\cal T}_{c}$, if each of its triangles and edges is combinatorially distinct. By this we mean that each triangle is defined uniquely by a set of three distinct vertices, and each edge by a set of two distinct vertices. This excludes e.g. the presence of two distinct edges that share the same vertices, or of two distinct triangles that have more than one edge in common. The graph dual to a 2D combinatorial triangulation is a trivalent $\phi^{3}$-graph without tadpole or self-energy subgraphs [35]. Yet another way to characterize elements of ${\cal T}_{c}$ uses the definitions just introduced: a triangulation is combinatorial if the link of each zero-simplex is homeomorphic to a one-sphere (cf. Fig. 14), and the link of each one-simplex is homeomorphic to a zero-sphere (the set of two points),